if log a =.250 and log b =.644, then what does log (b/a)=?
Answer by stanbon(75887)

if log a =.250 and log b =.644, then what does log (b/a)=?
---------
Log b/a = logb - loga = 0.644 - 0.250 = 0.394
===============
Cheers,
Stan H.